OpenStudy (anonymous):

given: f(x)=3x^2+4x and g(x)=2x^2-x+1 what is (f+g)(x) a. 6x^2-4x+1 b. 5x^2+3x+1 c. x^2+5x-1 d. -x^2-5x-1

geerky42 (geerky42):

Just combine like terms; \[(f+g)(x) = f(x) + g(x) = 3x^2 + 4x ~~~+~~~2x^2-x+1 \\~\\~~~~~~~~~~~~~~~~~= (3+2)x^2 + (4-1)x + 1 ~~=~~\boxed{5x^2 + 3x + 1} \]